# On the Eldan–Gross inequality

### v1

@@ Abstract.
[…]
@@ 1. Introduction
$|\nabla f|:=\left(\sum_{j=1}^{n}|D_{j}f|^{2}\right)^{1/2}.$
@@
$\mathbb{E}|\nabla f|\geq C\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sum_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)}.$ (2)
@@
$\mathbb{E}\sqrt{h_{f}}\geq C_{p}\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sum_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)},$ (3)
where $h_{f}$ is supported on a set $A=\{x:f(x)=1\}$, and at each point $x\in A$, the value $h_{f}$ counts the number of edges joining $x$ with $A^{c}$, i.e., complement of $A$. Clearly we always have $|\nabla f|=\sqrt{h_{f}+h_{-f}}\geq\sqrt{h_{f}}$, therefore, the conjecture (3), if correct, would imply Eldan– Gross inequality (2). However, in general $\mathbb{E}|\nabla f|$ and $\mathbb{E}\sqrt{h_{f}}$ are not comparable, for example, if $A=\{(1,\ldots,1)\}$ is a single point then $\mathbb{E}|\nabla f|=\frac{n+\sqrt{n}}{2^{n}}\gg\frac{\sqrt{n}}{2^{n}}=\mathbb{E}\sqrt{h_{f}}$.
The proof of Eldan and Gross of the inequality (2) uses stochastic analysis and also leads to other impressive consequences. There are now several other proofs of (2); see for example van Handel [Ros20], Eldan–Kindler–Lifshitz–Minzer [EKLM23] and Beltran–Ivanisvili–Madrid [BIM23, Remark 2]. The main aim of this paper is to illustrate that some key ingredients of these proofs extend to more general spaces and the Boolean cube structure is not essential. Our proof of Theorem 1 relies on hypercontractivity and an isoperimetric-type inequality by Bobkov and Götze [BG99]. We shall prove the Theorem 1 on the $p$-biased hypercube; see Section 2 for details.
@@
for suitable $f:E\to$. By “suitable" we refer to a certain class of functions for which $\Gamma(f)$ is well-defined. In the sequel, we omit the rigorous definition to keep presentation compact, and we refer to [BGL14, Chapter 3] for detailed definitions for $(E,\mu,\Gamma)$ being a full Markov triple. One crucial property here is the diffusion property, which may be understood as the chain rule for $L$, so that many functional inequalities follow from Bakry–Émery criterion
@@
For any $n\geq 1$, denote $E^{n}$ the $n$-fold product of $(E,\mu)$ equipped with the product probability measure $\mu_{n}.$ We consider the tensor product of $(P_{t})$ over $E^{n}$, which will be denoted by $(P_{t})=(e^{-tL})$ again whenever no confusion can occur. For $f:E^{n}\to$, we define its variance
@@
Here and in what follows, for $f:E^{n}\to$, the $L^{p}$ norm $\|f\|_{p}$ is defined with respect to $\mu_{n}$. We define
@@
$\mathbb{E}|\nabla f|\geq C_{K}^{\prime}\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sum_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)}.$ (6)
@@ Acknowledgment
P.I. is supported by National Science Foundation CAREER-DMS-2152401.
@@ 2. Boolean cube case
In this section, we give an alternative proof of Theorem 1. The proof also extends to the biased case which we now explain. For any $p\in(0,1)$ we equip $\{-1,1\}$ with the probability measure $\mu_{p}=p\delta_{1}+(1-p)\delta_{-1}$. We use $\|f\|_{q}$ to denote the $L^{q}$-norm of $f:\{-1,1\}^{n}\to$ with respect to $\mu_{p}^{\otimes n}$. We write $\mathbb{E}=\mathbb{E}_{\mu_{p}^{\otimes n}}$ for short, and consider
@@
which is a semigroup of unital positive linear operators over {${}^{\{-1,1\}^{n}}$ such that $P_{0}(f)=f$ and $\lim_{t\to\infty}P_{t}(f)=\mathbb{E}f$. Accordingly, we write $\textnormal{Var}(f)$ for $\textnormal{Var}_{\mu_{p}}(f)$ with respect to this biased $\mu_{p}$.
@@
and the constant factor of $4p(1-p)$ will make no essential difference for our discussion. So we keep our choice (8). We still put
$|\nabla f(x)|:=\left(\sum_{j=1}^{n}|D_{j}f(x)|^{2}\right)^{1/2}.$
@@
$\mathbb{E}|\nabla f|\geq C_{p}\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sum_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)}.$ (9)
One ingredient of the proof can be derived from an isoperimetric-type inequality of Bobkov and Götze [BG99]. The proof is inspired by the unbiased case [KOS08].
@@
$\mathcal{I}(\mathbb{E}_{\nu}h)\leq\sqrt{2}\mathbb{E}_{\nu}\sqrt{\sum_{i}\textnormal{Var}_{\nu_{i}}(h)}$ (12)
[…]
@@
$\varphi_{i}(g):=P_{t}(g)(x_{i})=(1-e^{-t})[pg(1)+(1-p)g(-1)]+e^{-t}g(x_{i}),\qquad g:\Omega_{i}=\{-1,1\}\to,$
@@
$\mathbb{E}_{\nu_{i}}(g)=\int_{\{-1,1\}}gd\nu_{i}=P_{t}(g)(x_{i}),\qquad g:\{-1,1\}\to.$
@@
where $a=h(y_{1},\dots,1,\dots,y_{n}),b=h(y_{1},\dots,-1,\dots,y_{n})$ and $c=h(y_{1},\dots,x_{i},\dots,y_{n})$. Here we also write $q=1-p$ and let us denote $r:=\max\{p,q\}$. Thus using $a=a^{2},b=b^{2},c=c^{2}$ we obtain by simple computations: if $c=a$, then
$\textnormal{Var}_{\nu_{i}}(h(y))=\mathbb{E}_{\nu_{i}}(h(y))-\mathbb{E}_{\nu_{i}}(f(y))^{2}=q(1-e^{-t})(p+qe^{-t})(a-b)^{2}\leq r^{2}(1-e^{-2t})(a-b)^{2},$ (14)
@@
$\textnormal{Var}_{\nu_{i}}(h(x))\leq r^{2}(1-e^{-2t})|h(x_{1},\dots,x_{i},\dots,x_{n})-h(x_{1},\dots,-x_{i},\dots,x_{n})|^{2}=4r^{2}(1-e^{-2t})|D_{j}h(x)|^{2}$ (16)
@@
$\sum_{i=1}^{n}\textnormal{Var}_{\nu_{i}}(h(x))\leq 4r^{2}(1-e^{-2t})\sum_{i=1}^{n}|D_{j}f(x)|^{2}=4r^{2}(1-e^{-2t})|\nabla h(x)|^{2}.$ (17)
@@
$1-(P_{t}f)^{2}\leq 2\max\{p,1-p\}\sqrt{\pi(1-e^{-2t})}P_{t}|\nabla f|,\qquad f:\{-1,1\}^{n}\to\{-1,1\}.$ (21)
@@
For $p\in(0,1)$ we consider the semigroup $P_{t}$ defined in (7). Then for any $21<r<$ 2 and for any $f:\{-1,1\}^{n}\to$ we have
@@
When $p=1/2$, this is a well-known result by Bonami [Bon70] and the time in (22) is optimal. In the general biased case, the optimal time is more involved [Ole03]. We give a proof of (22) in the Appendix for convenience.
From the above proposition, we may deduce the following lemma.
@@
Fix $p\in(0,1)$. For all $t\geq 0$ and for all $f:\{-1,1\}^{n}\to$, we have
$\textnormal{Var}(P_{t}f)\leq\left(\sum_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}\right)^{\theta(t)}\textnormal{Var}(f)^{1-\theta(t)},$ (23)
@@
$\textnormal{Var}(f_{t})=\sum_{j=1}^{n}\|f_{t}^{(j-1)}-f_{t}^{(j)}\|_{2}^{2}$ $=\sum_{j=1}^{n}\|P_{t}(f^{(j-1)}-f^{(j)})\|_{2}^{2}$
$\leq\sum_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1+e^{-2Kt}}^{2}$
$\leq\sum_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1}^{2\theta(t)}\|f^{(j-1)}-f^{(j)}\|_{2}^{2-2\theta(t)}$
$\leq\left(\sum_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1}^{2}\right)^{\theta(t)}\left(\sum_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{2}^{2}\right)^{1-\theta(t)}$
$=\left(\sum_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1}^{2}\right)^{\theta(t)}\textnormal{Var}(f)^{1-\theta(t)},$
@@
$\textnormal{Var}(f_{t})\leq\left(\sum_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}\right)^{\theta(t)}\textnormal{Var}(f)^{1-\theta(t)}.\qed$
@@
since $\textnormal{Var}h=\mathbb{E}h(1-\mathbb{E}h)$. Combining (26) and (28) finishes the proof. To see (27), it suffices to consider the cases when $x\downarrow 0$ and $x\uparrow 1$. Note that $\mathcal{I}(x)=\mathcal{I}(1-x)$, so it suffices to show (27) for $x\downarrow 0$ and this is known [Bob97, Remark 1]. ∎
@@
Denote $W(f):=\sum_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}$. Recalling (11), the desired inequality (9) follows immediately when $W(f)$ is large enough, say $W(f)\geq 1/100$. So let us assume $W(f)<1/100$.
@@ 3. Continuous analogues
    Hypercontractivity [BGL14, Theorem 5.2.3, Proposition 5.7.1]: for $f:E\to$, $t\geq 0$ and $1<p\leq q<\infty$ we have
@@
    (Local) Bobkov’s inequality [BGL14, Theorem 8.5.3]: for $g:E\to\{0,1\}$ and $t\geq 0$ we have under suitable approximation
@@
Here, (36) is a special case of [BGL14, Theorem 8.5.3] that suffices our use. For $n\geq 1$, we shall still denote by $P_{t}$ its $n$-fold products on $E^{n}$. Then above results actually hold for $E^{n}$ after a tensorization argument, which we will use directly.
@@
Under the above assumptions, for $f:E^{n}\to$, we have
$\textnormal{Var}(P_{t}f)\leq\left(\sum_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}\right)^{\theta_{K}(t)}\textnormal{Var}(f)^{1-\theta_{K}(t)},$ (39)

### v2

@@ Abstract.
$\intop\nolimits_{\{-1,1\}^{n}}\sqrt{s_{f}(x)}d\mu(x)\geq C\textnormal{Var}(f)\sqrt{\log\left(1+\frac{1}{\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)}$
@@ 1. Introduction
$|\nabla f|:=\left(\sumop\displaylimits_{j=1}^{n}|D_{j}f|^{2}\right)^{1/2}.$
@@
$\mathbb{E}|\nabla f|\geq C\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)}.$ (2)
@@
$\mathbb{E}\sqrt{h_{f}}\geq C_{p}\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)},$ (3)
where $h_{f}$ is supported on a set $A=\{x:f(x)=1\}$, and at each point $x\in A$, the value $h_{f}$ counts the number of edges joining $x$ with $A^{c}$, i.e., complement of $A$. Clearly we always have $|\nabla f|=\sqrt{h_{f}+h_{-f}}\geq\sqrt{h_{f}}$, therefore, the conjecture (3), if correct, would imply Eldan–Gross inequality (2). However, in general $\mathbb{E}|\nabla f|$ and $\mathbb{E}\sqrt{h_{f}}$ may behave quite differently as $n\to\infty$, for example, if $A=\{(1,\ldots,1)\}$ is a single point then $\mathbb{E}|\nabla f|=\frac{n+\sqrt{n}}{2^{n}}\gg\frac{\sqrt{n}}{2^{n}}=\mathbb{E}\sqrt{h_{f}}$.
The proof of Eldan and Gross of the inequality (2) uses stochastic analysis and also leads to other impressive consequences. There are now several other proofs of (2); see for example van Handel [Ros20], Eldan–Kindler–Lifshitz–Minzer [EKLM23] and Beltran–Ivanisvili–Madrid [BIM23, Remark 2]. The main aim of this paper is to illustrate that some key ingredients of these proofs extend to more general spaces and the Boolean cube structure is not essential. Let us remark that the proof in [Ros20] can achieve the same goal after some modifications. Here, we highlight the use of an isoperimetric-type inequality by Bobkov and Götze [BG99] in 1999, the way of using which can be of independent interest.
We shall prove Theorem 1 on the $p$-biased hypercubes; see Section 2 for details. The proof relies on Bobkov and Götze’s isoperimetric-type inequality [BG99] and the hypercontractivity.
@@
for suitable $f:E\to\real$. By “suitable" we refer to a certain class of functions for which $\Gamma(f)$ is well-defined. In the sequel, we omit the rigorous definition to keep presentation compact, and we refer to [BGL14, Chapter 3] for detailed definitions for $(E,\mu,\Gamma)$ being a full Markov triple. One crucial property here is the diffusion property, which may be understood as the chain rule for $L$, so that many functional inequalities follow from Bakry–Émery criterion
@@
For any $n\geq 1$, denote $E^{n}$ the $n$-fold product of $(E,\mu)$ equipped with the product probability measure $\mu_{n}.$ We consider the tensor product of $(P_{t})$ over $E^{n}$, which will be denoted by $(P_{t})=(e^{-tL})$ again whenever no confusion can occur. For $f:E^{n}\to\real$, we define its variance
@@
Here and in what follows, for $f:E^{n}\to\real$, the $L^{p}$ norm $\|f\|_{p}$ is defined with respect to $\mu_{n}$. We define
@@
$\mathbb{E}|\nabla f|\geq C_{K}^{\prime}\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)}.$ (6)
@@ Acknowledgment
P.I. is supported by National Science Foundation CAREER-DMS-2152401. H.Z. is supported by National Science Foundation DMS-2453408.
@@ 2. Boolean cube case
In this section, we give an alternative proof of Theorem 1. The proof also extends to the biased case which we now explain. For any $p\in(0,1)$ we equip $\{-1,1\}$ with the probability measure $\mu_{p}=p\delta_{1}+(1-p)\delta_{-1}$. We use $\|f\|_{q}$ to denote the $L^{q}$-norm of $f:\{-1,1\}^{n}\to\real$ with respect to $\mu_{p}^{\otimes n}$. We write $\mathbb{E}=\mathbb{E}_{\mu_{p}^{\otimes n}}$ for short, and consider
@@
which is a semigroup of unital positive linear operators over $\real{}^{\{-1,1\}^{n}}$ such that $P_{0}(f)=f$ and $\lim_{t\to\infty}P_{t}(f)=\mathbb{E}f$. Accordingly, we write $\textnormal{Var}(f)$ for $\textnormal{Var}_{\mu_{p}}(f)$ with respect to this biased $\mu_{p}$.
@@
and the constant factor of 4p(1p)4p(1-p) will make no essential difference for our discussion. So we keep our choice (8). We still put
$|\nabla f(x)|:=\left(\sumop\displaylimits_{j=1}^{n}|D_{j}f(x)|^{2}\right)^{1/2}.$
@@
$\mathbb{E}|\nabla f|\geq C_{p}\textnormal{Var}(f)\sqrt{\log\left(1+\frac{e}{\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}}\right)}.$ (9)
One ingredient of the proof can be derived from an isoperimetric-type inequality of Bobkov and Götze [BG99]. The proof is inspired by the unbiased case [KOS08]. More discussions can be found in Appendix B. Similar estimates were obtained in [Ros20, Claim 8, items 1-2] via a more direct argument.
@@
$\mathcal{I}(\mathbb{E}_{\nu}h)\leq\sqrt{2}\mathbb{E}_{\nu}\sqrt{\sumop\displaylimits_{i}\textnormal{Var}_{\nu_{i}}(h)}$ (12)
for any measurable $h:\Omega\to\{0,1\}$, where $\mathcal{I}$ is the Gaussian isoperimetric profile $\mathcal{I}:[0,1]\to[0,1/\sqrt{2\pi}]$ given by $\mathcal{I}=\varphi\circ\Phi^{-1}$ with $\Phi(r)=\intop\nolimits_{-\infty}^{r}\varphi(x)dx$ and $\varphi(x)=\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}$.
@@
$\varphi_{i}(g):=P_{t}(g)(x_{i})=(1-e^{-t})[pg(1)+(1-p)g(-1)]+e^{-t}g(x_{i}),\qquad g:\Omega_{i}=\{-1,1\}\to\real,$
@@
$\mathbb{E}_{\nu_{i}}(g)=\intop\nolimits_{\{-1,1\}}gd\nu_{i}=P_{t}(g)(x_{i}),\qquad g:\{-1,1\}\to\real.$
@@
where $a=h(y_{1},\dots,1,\dots,y_{n}),b=h(y_{1},\dots,-1,\dots,y_{n})$ and $c=h(y_{1},\dots,x_{i},\dots,y_{n})$. Here we also write q=1pq=1-p and let us denote $r:=\max\{p,q\}$. Thus using $a=a^{2},b=b^{2},c=c^{2}$ we obtain by simple computations: if $c=a$, then
$\textnormal{Var}_{\nu_{i}}(h(y))=\mathbb{E}_{\nu_{i}}(h(y))-\mathbb{E}_{\nu_{i}}(h(y))^{2}=q(1-e^{-t})(p+qe^{-t})(a-b)^{2}\leq r^{2}(1-e^{-2t})(a-b)^{2},$ (14)
@@
$\textnormal{Var}_{\nu_{i}}(h(x))\leq r^{2}(1-e^{-2t})|h(x_{1},\dots,x_{i},\dots,x_{n})-h(x_{1},\dots,-x_{i},\dots,x_{n})|^{2}=4r^{2}(1-e^{-2t})|D_{i}h(x)|^{2}$ (16)
@@
$\sumop\displaylimits_{i=1}^{n}\textnormal{Var}_{\nu_{i}}(h(x))\leq 4r^{2}(1-e^{-2t})\sumop\displaylimits_{i=1}^{n}|D_{i}f(x)|^{2}=4r^{2}(1-e^{-2t})|\nabla h(x)|^{2}.$ (17)
@@
$1-(P_{t}f)^{2}\leq 2\max\{p,1-p\}\sqrt{\pi(1-e^{-2t})}P_{t}|\nabla f|,\qquad f:\{-1,1\}^{n}\to\{-1,1\}.$
@@
For $p\in(0,1)$ we consider the semigroup $P_{t}$ defined in (7). Then for any $1<r<2$ and for any $f:\{-1,1\}^{n}\to\real$ we have
@@
When $p=1/2$, this is a well-known result by Bonami [Bon70] and the time in (21) is optimal. In the general biased case, the optimal time is more involved [Ole03]. We give a proof of (21) in the Appendix A for convenience.
From the above proposition, one may deduce the following lemma whose proof is essentially the same as that of [Ros20, Lemma 11].
@@
Fix $p\in(0,1)$. For all $t\geq 0$ and for all $f:\{-1,1\}^{n}\to\real$, we have
$\textnormal{Var}(P_{t}f)\leq\left(\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}\right)^{\theta(t)}\textnormal{Var}(f)^{1-\theta(t)},$ (22)
@@
$\textnormal{Var}(f_{t})=\sumop\displaylimits_{j=1}^{n}\|f_{t}^{(j-1)}-f_{t}^{(j)}\|_{2}^{2}$ $=\sumop\displaylimits_{j=1}^{n}\|P_{t}(f^{(j-1)}-f^{(j)})\|_{2}^{2}$
$\leq\sumop\displaylimits_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1+e^{-2Kt}}^{2}$
$\leq\sumop\displaylimits_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1}^{2\theta(t)}\|f^{(j-1)}-f^{(j)}\|_{2}^{2-2\theta(t)}$
$\leq\left(\sumop\displaylimits_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1}^{2}\right)^{\theta(t)}\left(\sumop\displaylimits_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{2}^{2}\right)^{1-\theta(t)}$
$=\left(\sumop\displaylimits_{j=1}^{n}\|f^{(j-1)}-f^{(j)}\|_{1}^{2}\right)^{\theta(t)}\textnormal{Var}(f)^{1-\theta(t)},$
@@
$\textnormal{Var}(f_{t})\leq\left(\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}\right)^{\theta(t)}\textnormal{Var}(f)^{1-\theta(t)}.\qed$
@@
since $\textnormal{Var}(h)=\mathbb{E}h(1-\mathbb{E}h)$. Combining (25) and (27) finishes the proof. To see (26), it suffices to consider the cases when $x\downarrow 0$ and $x\uparrow 1$. Note that $\mathcal{I}(x)=\mathcal{I}(1-x)$, so it suffices to show (26) for $x\downarrow 0$ and this is known [Bob97, Remark 1]. ∎
@@
Denote $W(f):=\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}$. Recalling (11), the desired inequality (9) follows immediately when $W(f)$ is large enough, say $W(f)\geq 1/100$. So let us assume $W(f)<1/100$.
@@ 3. Continuous analogues
    Hypercontractivity [BGL14, Theorem 5.2.3, Proposition 5.7.1]: for $f:E\to\real$, $t\geq 0$ and $1<p\leq q<\infty$ we have
@@
    (Local) Bobkov inequality [BGL14, Theorem 8.5.3]: for $g:E\to\{0,1\}$ and $t\geq 0$ we have under suitable approximation
@@
Here, (35) is a special case of [BGL14, Theorem 8.5.3] that suffices for our use. In fact, the original [BGL14, Theorem 8.5.3] states that for $\alpha\geq 0,t\geq 0$ and measuable functions $f:E\to\real$
$\sqrt{\mathcal{I}^{2}(P_{t}f)+\alpha\Gamma(P_{t}f)}\leq P_{t}\left(\sqrt{\mathcal{I}^{2}(f)+c_{\alpha}(t)\Gamma(f)}\right),\qquad c_{\alpha}(t)=\frac{1-e^{-2Kt}}{K}+\alpha e^{-2Kt}.$ (36)
Then (35) follows by taking $\alpha=0$ in (36) and approximating the characteristic function $g:E\to\{0,1\}$ in (35) with measurable functions $f$ in (36). Here, the Boolean property plays a role, and we used $\mathcal{I}(0)=\mathcal{I}(1)=0$. For $n\geq 1$, we shall still denote by $P_{t}$ its $n$-fold products on $E^{n}$. Then the above results actually hold for $E^{n}$ after a tensorization argument, which we will use directly.
@@
Under the above assumptions, for $f:E^{n}\to\real$, we have
$\textnormal{Var}(P_{t}f)\leq\left(\sumop\displaylimits_{j=1}^{n}\textnormal{Inf}_{j}(f)^{2}\right)^{\theta_{K}(t)}\textnormal{Var}(f)^{1-\theta_{K}(t)},$ (39)